# Human-AI Collaboration: Trade-offs Between Performance and Preferences

### v1

@@ Abstract
[…]
1 Main
Contemporary AI technologies have matured to the point where their integration into everyday activities has become feasible. This integration is taking place in a wide range of fields including healthcare, education, and gaming [1]. In many of these settings, the user explicitly instructs the AI to produce an output or the AI simply offers non-binding suggestions to the user. While this paradigm in which the AI assists the human remains dominant, there is growing interest in collaborative frameworks where both humans and AI make independent contributions to shared goals. Unlike AI-assisted decision-making tasks where AI supports human decision-makers [2], such human-AI collaboration involves both the AI and the human taking actions that directly influence outcomes, whether in real-world scenarios or simulated environments [3, 4, 5, 6, 7, 8, 9]. Despite significant progress in multi-agent AI research, much of the current work excludes humans from the research design, potentially leading to AI systems that are not well aligned with human goals and preferences [10]. This study aims to build effective, human-aligned, collaborative AI by combining performance-driven AI designs and human-centered design studies. Specifically, we evaluate the impact of distinct AI collaboration strategies on performance and human users perceptions. Ultimately, this study seeks to formalize a critical trade-off between an AI’s performance and its acceptance by human users.
@@ 1 Main
Current machine learning research typically views collaborative AI as a multi-agent reinforcement learning (MARL) problem [11]. Although MARL algorithms can perform well on benchmark metrics, they often overlook human-computer interaction (HCI) design principles [3]. This divergence is concerning, as the adoption of these algorithms will critically depend on human users accepting these AI agents. The focus on performance metrics over user studies is in part due to the complexity of modeling human behavior. This is not surprising, given the nuance in human behavior and the lack of formal overarching models that could anticipate these intricacies [12, 3, 5, 10]. Due to this challenge of modeling nuanced human behaviors, research into human-AI collaboration generally focuses on either multi-agent performance metrics or human-centered designs, rarely both [12, 5]. Attempts to integrate these important human considerations often add complexity to the modeling framework. For example, researchers have modeled agents from human demonstrations (behavior cloning). This method has several notable limitations. Firstly, human data is relatively costly to collect. Secondly, training an algorithm to reproduce human-like behavior does not explicitly integrate validated design principles [13]. Finally, behavior-cloning struggles to perform as well as more simulation-based methods [6]. This combination of circumstances limits behavior-cloning to niche situations, as it is neither performance-maximizing, nor interpretable.
@@
Subsequently, the state of human-in-the-loop multi-agent systems effectively neglects recent research, which has shown that effective collaborative AI must take into account subjective factors beyond objective performance measures [4, 14, 5, 15, 10, 16, 17, 18]. For example, people prefer an AI agent whose behavior is predictable and transparent, as these characteristics make the AI’s actions more understandable and reliable [4, 14, 16, 17, 18]. Similarly, people prefer non-adaptive, rule-based agents over learning-based agents due to their predictability and ease of interaction [10]. This is in stark contrast to state-of-the-art methods in MARL, which involve complex and adaptive models that are inherently opaque in their decision-making [11]. As such, there is a definite need for paradigms that can help close this widening gap.
There is a body of research that should inform AI development, but remains challenging to integrate in practice. Recent research in human-computer interaction has identified a number of behavioral characteristics, or “traits”, that contribute to successful human-AI collaboration; including perceptions of warmth, competence, intentionality, and fun [5, 10]. These features approximately map onto the collaborator’s ability to relate to the human in a considerate, communicative, and engaging way. Despite these findings, most research stops short of explicitly designing agents with behavioral characteristics that humans find desirable in collaborative tasks [10]. There are only few exceptions, which leverage preexisting theories that lend themselves to algorithmic implementations
@@
In this study, we seek to address the gap between performance-oriented algorithm design and user-centered preferences in human-AI collaboration. To do this, we developed a set of rule-based agents, each representing variations of an egocentric, performance-maximizing agent that incorporates additional manipulations meant to reflect one or more of these behavioral traits. In addition, we conducted a behavioral study that systematically evaluates how these algorithmic variations impact both the performance of the human-AI team and the perceived traits of the AI from the human’s perspective.
@@
Our aim is to address two central questions. First, which factors most influence human preferences for collaborative AI agents? In our experiments, we examined both objective metrics related to performance as well as subjective factors related to people’s perception of the AI agents’ various behavioral traits. Second, do the trade-offs between AI performance and human preference always operate in a strictly linear, zero-sum manner, where improving one inherently detracts from the other? Alternatively, are there strategies or design choices that can achieve a net positive effect—where a marginal decrease in AI performance (if any) is more than compensated by significant improvements in human approval?
@@
1.1 Behavioral Experiments
This study presents an empirical framework to explore how different aspects of agent behavior lead to effective human-AI collaboration. Through behavioral experiments, we conducted a systematic evaluation of various AI agent designs, examining how different collaborative strategies affect both performance and human preferences. In our study, human participants interacted with several variations of a collaborative AI agent within a dynamic decision-making task. By analyzing participants’ experiences and preferences when collaborating with different types of agents, we seek to identify the key factors that contribute to AI collaboration. Importantly, we provide a paradigm in which different algorithmic manipulations can be mapped onto subjective perceptions, allowing us to evaluate whether our manipulations approximate commonly reported design principles. Thus, our experiments provide an approach for bridging the gap between performance-oriented algorithm design and user-centered preferences in human-AI collaboration.
1.1.1 Collaborative Target Interception Task
To investigate human preferences for collaborative AI agents, we created a task in which a human player and an AI agent work together towards a common goal: achieving a maximally high team score. Our task is an extension of a dynamic decision-making task previously used to study how humans adopt AI assistance [19]. The objective of the game is to collect as many points as possible by intercepting point-valued targets which move at constant speeds through the game environment (see Figure 1). The task has some of the planning requirements of traveling salesman problems that have been studied in the context of human problem solving [20], although it additionally involves moving destinations. Importantly, the task necessitates collaborative planning between the human player and the AI agent, as targets vanish after being intercepted by a player or exiting the game-view. New targets appear at random intervals, meaning inconsiderate collaborators can ultimately get in each other’s way. Each player clicks on targets to direct their avatar to the best interception point. This means that the interface handles the navigation while player’s focus on the decision-making.
@@ 1.1.2 Collaborative Agents
2 Results
@@
We begin our reporting of the results by showing outcomes from objective performance metrics such as the performance of human and collaborative AI agents. Additionally, we highlight behavioral measures related to the proximity between the agents and the degree to which one agent interferes with the plan of the other agent. We then report the results from the subjective metrics based on the questionnaire responses. Finally, we examine human preferences for various types of collaborative agents and apply predictive models to determine which objective and subjective metrics best predict choice.
2.1 Objective Metrics
2.1.1 Performance Differences
@@
Figure 3 shows the performance of the human-AI team where the score is combined across the human player and the AI agent. In the high target density condition, human+Omit teams slightly outperformed the human+Ignorant teams, although not statistically significant, $BF_{10}<1$. Thus, the additional design features of the Omit agent caused human performance gains that were at least equally as high than the AI performance decreases. In the low target density condition, the best team performance was achieved with the Ignorant agent closely followed by the Omit agent.
@@
Figure 2: Performance of the human and AI player by AI agent type (columns) and target density (rows). Performance is assessed by a relative score: the total points scored relative to the total points that were available during game play. Gray areas visualize the distribution of proportional scores; error bars show the standard error of the mean.
@@
Figure 3: Mean team score by AI agent type and target density. The team score is based on the sum of score of the AI agent and the human playing with that AI agent relative to the total value of points that was available during game play. Gray shading indicates the distribution of values, while error bars show the standard error from the mean.
2.1.2 Other Behavioral Metrics
@@
We also examined several behavioral metrics that distinguish between the AI agents, focusing on measures related to conflict between the human player and the AI agent. One such metric, the number of AI ’steals’, is defined as the number of times the AI agent intercepts a target initially pursued by the human player. Appendix Figure B2 presents a visualization of these results. As expected, the Ignorant agent shows the highest number of interceptions of targets that the human intended to catch, since this agent disregards the human’s intentions and will pursue targets regardless of the human’s planned actions.
2.2 Subjective Metrics: Questionnaire Responses
@@
Figure 4: Mean questionnaire scores by AI agent type and target density. Questions were rated on a 7-point Likert-scale. Q1: “The bot and I were a team”, Q2: “The bot was competent”, Q3: “I understood the bot’s intentions”, Q4: “The bot understood my intentions”, Q5: “I contributed more to the team’s performance”, Q6: “The bot was easy to play with”, Q7: “The bot was fun to play with”, Q8: “The bot had a similar playing style to me”. Error bars indicate the standard error from the mean.
@@
Figure 4 shows the questionnaire results. The most general finding that holds for all items except Q3 (“I understood the bot’s intentions”) is that there are significant differences in the ratings for the Ignorant agent, compared to all other agents. These statistical findings reveal a difference between the Ignorant agent and the other agents. However, for some items this pattern is more nuanced as demonstrated by the significant target density interaction effects observed for items Q2, Q5, Q8, $BF_{10}>100$. Exceptions include Q2 (“The bot was competent”) where in the low density condition the Ignorant, Divide, and Delay agents were rated equally well, while in the high density condition the Bottom-feeder was rated worse than all other agents, including the Ignorant agent. Furthermore, comparisons in the high density condition of Q8 (“The bot had a similar playing style to me”) show that the ratings of the Ignorant and Bottom-Feeder agents were roughly equally low, while all other agents received significantly higher ratings. Q3 evinced no differences in ratings across agents.
@@
Table 1 shows example responses when participants were asked to explain their choice. Participants voice many of the human-centered design considerations in their open-ended responses. The theme of teaming was highly represented in our participants’ open-ended responses. Participants frequently pointed out that the Ignorant agent was not being a good teammate. Appendix 7.4 provides a content analysis that confirms that the majority of open-ended responses focused on teaming.
Table 1: Examples of explanations provided in the open-ended surveys.
@@
Teaming “The copper bot felt less like competition and more like a fellow teammate. When I would choose a target, even if it was originally planning on going to that target, it would get out of the way and let me grab the target. This seemed more aligned with two teammates working together than the other robot.” “We worked as a team compared to the other bot” “It felt like we were a team and just trying to collect as many circles as possible where as the copper bot felt like it was competing against me and would go change direction based on the highest point circles rather than holding down an area of the platform like me and the blue bot would.”
Likability “The blue bot made playing the game easy and fun. The blue bot spent most of its time in a quadrant away from where I was playing, allowing me not to feel crowded or pressured.”
Intentionality “The Purple bot seemed to just let me get any of the targets that I wanted and didn’t try to fight for them.”
2.3 Preferences for Collaborative Agents
@@
Figure 5: Choice preferences across pairs of agents for each target density condition. Each matrix cell indicates the percentage of participants preferring the row-associated agent A over the column-associated agent B. For instance, 82% of participants preferred the Omit agent over the Ignorant agent in the low target density condition. The side column shows the overall preference percentage for each row agent across all pairings. Asterisks denote choice percentages that significantly deviate from 50%, indicated by a Bayes factor greater than 10. Note that results here are averaged across different presentation orders of the agents (e.g., agent A could have been presented first or second in the experiment).
2.4 Predictive Models for Human Preferences
@@
To estimate the weights in the logistic regression model of Equation 2, we use Bayesian methods. We separately apply the model to objective and subjective metrics, further breaking down the results by target density. For the objective metrics, we included features such as the human and AI scores, score inequality (defined as the absolute difference between human and AI scores), the number of AI steals, and the number of path intersections between the human and AI agent. For subjective metrics, we included Likert ratings from each of the 8 survey questions. Figure 6 presents the posterior estimates for the $\beta$ weights (see Appendix Table D1 for full results). A positive $\beta$ weight indicates that a larger positive difference in feature values between Agents $X$ and $Y$ increases the likelihood of choosing Agent $X$.
@@
Another key predictor is AI "steals", with agents scoring higher on these features being less preferred, although there is only evidence for this effect in the low target density condition ($\emph{BF}>100$), where there is more opportunity for competitive interaction.
@@ 2.4.1 Predictive Accuracy
Figure 6: Posterior ($\beta$) coefficients of Bayesian logistic regression models predicting choice. The coefficients are shown across objective and subjective metrics (left and right panels) and target densities (indicated by colors). Coefficient estimates can be thought of as weights for the importance of metrics in explaining choices. Error bars represent 95% credible intervals.
Figure 7: Scatter plot of mean relative AI agent score versus the human score for each type of human-AI team, separated by target density. Color is reflective of the probability of the AI type being preferred in pairwise match-ups. The diagonal represents the points of equal scores for the AI agent and the human player. The probability with which an agent is chosen appears inversely related to the distance from the diagonal, suggesting that human players have a preference for score equality. Error bars are indicative of the standard error from the mean.
Figure 8: Scatter plot of team performance over probability of being chosen, by AI agent condition and target density conditions. Error bars reflect the standard error from the mean.
2.4.2 Trade-off Between Performance and Preference

### v2

@@ Abstract
keywords:
Human-AI collaboration, Dynamic Decision-Making, Human-Centered Algorithm Design, Bayesian Modeling of Preferences
Significance
Human-AI collaboration is expected to grow in the coming years. Particular attention is being paid to agentic cooperative AI that is capable of autonomously performing helpful tasks without repeated human instruction due to its potential to significantly improve the performance of human-AI teams. However, the use of cooperative AI agents poses two key challenges: (1) the development of such agents in modern multi-agent reinforcement learning paradigms often excludes human collaborators, and (2) the process of integrating human preferences into the algorithms underlying AI agents remains poorly understood. Our study addresses these shortcomings by establishing an empirical framework to evaluate how algorithmic changes can be mapped to human preferences. Our study reveals key dynamics, such as algorithm changes that increase human liking of the AI agent without harming the performance of the human-AI team, and a pronounced human preference for inequity-aversion. These findings inform human-AI development by demonstrating how collaborative AI can be both effective and enjoyable. Our approach adjusts agent behavior by modifying algorithmic inputs and outputs, making it broadly applicable to new and existing agentic systems.
Introduction
Contemporary AI technologies have matured to the point where their integration into everyday activities has become feasible. This integration is taking place in a wide range of fields including healthcare, education, and gaming [19]. One setting for AI integration that is becoming increasingly common involves a user prompting an AI, for example a chatbot. Here, the user explicitly instructs the AI to produce an output or the AI simply offers non-binding suggestions to the user [2, 33]. In these types of interactions human-AI collaboration happens sequentially: A user prompt is followed by an AI response with the human always remaining the ultimate decision-maker. One alternative setting that has gained interest recently is collaboration with an agentic AI, where an AI agent can take actions independently from the human [6, 9, 20, 30, 3, 25, 22]. An AI being able to act independently from the human could purportedly yield productivity gains due to the ability to concurrently distribute labor across both the human and this agentic AI. For example, imagine that a human and an AI are collaborating on a software project with multiple outstanding tasks. Instead of merely advising the human on how to tackle each task, as in the chatbot setting, the agent might independently address some of the more routine tasks, thus leaving more time for the human to focus on the more challenging tasks.
@@ Introduction
Current machine learning research typically views agentic AI designed to collaborate with people as a multi-agent reinforcement learning (MARL) problem [11]. A common difficulty in MARL research is integrating human considerations into the algorithms that give rise to the agentic AI’s behavior. For example, researchers have modeled agents from human demonstrations (behavior cloning), but this method has several notable limitations. First, human data is relatively costly to collect. Second, training an algorithm to reproduce human-like behavior does not explicitly integrate validated design principles [8]. Finally, behavior cloning struggles to perform as well as more simulation-based methods [30]. This combination of factors limits the applicability of behavior-cloning approaches.
@@
Due to the complexity of integrating humans in the modeling, the development of human-in-the-loop multi-agent systems often neglects recent behavioral studies, which have shown that effective collaborative AI should take into account subjective factors beyond objective performance measures [9, 12, 20, 24, 28, 31, 37, 36]. For example, people prefer an AI agent whose behavior is predictable and transparent, as these characteristics make the AI’s actions more understandable and reliable [9, 12, 31, 37, 36]. Similarly, people prefer non-adaptive, rule-based agents over learning-based agents due to their predictability and ease of interaction [28]. The contradiction is evident when we compare human preferences like predictability and simplicity to a contemporary MARL algorithm, which involves complex and adaptive behaviors that are inherently opaque in their decision-making [11]. Finally, there is evidence to suggest that people expect collaborative AI agents to exhibit certain behavioral characteristics, or “traits”. These traits include behaviors that elicit perceptions of warmth, competence, intentionality, and fun [20, 28]; concepts that receive little attention in algorithm development. Given this divide between the technical development of agentic AI systems and people’s expectations for a collaborative agent, there is a growing need for paradigms that can shape algorithm development in a manner that is compatible with human preferences.
@@
Addressing human preferences when developing collaborative AI agents is of significant concern since the adoption of AI agents will critically depend on human users’ acceptance [29], yet research concerned with the development of collaborative AI is generally focused on either multi-agent performance metrics or human-centered designs, rarely both [4, 20]. In this study, we aim to show an avenue for building effective, human-aligned, collaborative AI by combining performance-driven AI designs and human-centered design approaches. Specifically, we evaluate the impact of distinct AI collaboration strategies on performance and human users’ perceptions. To do this, we developed a set of rule-based agents, each representing variations of an egocentric, performance-maximizing agent that incorporates additional manipulations meant to reflect one or more behavioral traits. In addition, we conducted two behavioral experiments that systematically evaluate how these algorithmic variations impact both the performance of the human-AI team and the perceived traits of the AI from the human’s perspective.
@@
Our study addresses two central questions. First, which factors most influence human preferences for collaborative AI agents? In our experiments, we examined both objective metrics related to performance as well as subjective factors related to people’s perception of the AI agents’ various behavioral traits. Second, do the trade-offs between AI performance and human preference always operate in a strictly linear zero-sum manner where improving one inherently detracts from the other? Alternatively, are there strategies or design choices that can achieve a net positive effect—where a marginal decrease in AI performance (if any) is more than compensated by significant improvements in human performance and approval?
@@
Behavioral Experiments
To address these questions, we conducted two behavioral experiments to systematically evaluate various AI agent designs, examining how different collaborative strategies affect both performance and human preferences. In our study, human participants interacted with several variations of a collaborative AI agent within a dynamic decision-making task. By analyzing participants’ experiences and preferences when collaborating with different types of agents, we seek to identify the key factors that contribute to effective AI collaboration. Importantly, we provide a paradigm in which different algorithmic manipulations can be mapped onto subjective perceptions, allowing us to evaluate whether our manipulations approximate commonly reported design principles. Thus, our experiments provide an approach for bridging the gap between performance-oriented algorithm design and user-centered preferences in human-AI collaboration.
Collaborative Target Interception Task
To investigate human preferences for collaborative AI agents, we created a task in which a human player and an AI agent work together towards a common goal: achieving a maximally high team score. Our task is an extension of a dynamic decision-making task previously used to study how humans adopt AI assistance [16]. The objective of the game is to collect as many points as possible by intercepting point-valued targets which move at constant speeds through the game environment (see Figure 1). The task has some of the planning requirements of traveling salesman problems that have been studied in the context of human problem solving [10], although it additionally involves moving destinations. Importantly, the task necessitates collaborative planning between the human player and the AI agent, as targets vanish after being intercepted by a player or exiting the game view. New targets appear at random intervals, meaning inconsiderate collaborators can ultimately get in each other’s way. Each player clicks on targets to direct their avatar to the best interception point. This means that the interface handles the navigation while player’s focus on the decision-making.
@@
Experiment 1
Methods
Participants
Game Environment
There are two agents in the game: a collaborative AI player and a human player, each of whom have their own unique icon and distinctively colored square. The objective of the game is to intercept as many moving, point-valued targets as possible within a fixed time frame. Both the human player and the AI player can intercept targets but each target can be intercepted only once. Optimal task performance necessitates quick strategic decisions from the human player to intercept targets in a particular sequence during the limited time they are available while also paying attention to and coordinating with the actions of the collaborative AI player.
Targets spawn randomly at the edge of the circular game area. Their initial movement angle is randomly set within a cone of possible angles. Spawned targets move in straight-line directions at constant speed, sampled from a uniform distribution that ensures target speeds are between 1-50% slower than the player’s avatar. Targets exit the playable area if they are not intercepted. The spawning process ensures that the number of objects in the game area is limited to the target density (either 5 or 15 targets). One key feature of the spawning process is that it is independent of the player’s skill in interception. After a player intercepts a target, it disappears from the game area, but its path is still computed until it hits the perimeter. Only once a target, visible or not, hits this perimeter will a new target be spawned. Each target is worth between zero and fifteen points, with the probability distribution of point values following a Beta(1,2) distribution that we discretized over point values via binning. In practice, this means that low-value targets appear more often than high-value targets.
Collaborative AI Agents
Search Algorithm
The search algorithm is designed to approximate optimal solutions to the target interception task. The algorithm computes all possible interception sequences involving up to three targets, updating the positions of both the AI player and the targets throughout the sequence. This ensures that the AI player can respond to dynamic changes in the game state. The three-target limit is imposed to ensure the algorithm can operate in real time during behavioral experiments. For more details, see Appendix B.
Agent Variations
The search algorithm formed the basis for all agents in our experiment. We developed several variations of the search algorithm to create different collaborative AI players. These variations included changes to the target consideration set (which targets the AI player could pursue), delays in initiating a plan, and perception of point values. Variations were conceived as mechanisms that incorporate and give rise to heightened perceptions of traits observed in previous research [20, 28]. For a graphical overview of agent types see Figure 2. This study is not designed to test all possible combinations of these features. Instead, we focused on a set of five agents that test out a key set of variations:
  1. 1.
  2. 2.
  3. 3.
    Divide Agent: This agent operates in the same fashion as the omit agent but applies a divide and conquer strategy. This was done to make it easier for the human player and AI agent to avoid getting in each other’s way and, potentially, have better task delegation [35, 3]. The agent (virtually) divides the game area into two halves where the dividing line is orthogonal to the imaginary line from the human player to the game’s center-point. The agent only considers targets that can be intercepted in the half not occupied by the human player, with the allotted area being continuously recomputed as the human’s position changes. Therefore, this agent omits a larger set of targets from consideration than the omit agent, leaving more targets for the human player, further reducing potential interception conflicts between the players.
  4. 4.
  5. 5.
Figure 2: Overview of AI collaborator behavior differences. a. The Ignorant agent always pursues the highest value target, no matter what the human does. b. The Omit agent “omits” targets that the human is intended or predicted to intercept from consideration and is equivalent to Ignorant otherwise. c. The Divide agent extends the logic of Omit by also only considering targets on its half of the game environment. d. The Delay agent approximates the reaction time the human is demonstrating and is otherwise equivalent to Omit. e. The Bottom-Feeder “inverts” the value function of Omit, so it always pursues the lowest value target. Note that grayed out targets are not visible to the search algorithm.
Procedure
Figure 3: Illustration of the procedure in Experiment 1. Top and bottom rows (1 and 2) illustrate the two blocks in the experiment. Within each block, participants play two rounds, one with each AI agent from their assigned pair. Agents are then evaluated on a variety of dimensions using 7-point Likert scales. After submitting their ratings, participants indicate their preferred agent in a two-alternative forced choice. Finally, participants are asked to provide free-text responses explaining why they chose the agent they preferred. This procedure is repeated over two blocks where target density is varied. In the illustration, the first and second blocks have low and high target densities respectively. The density order is counterbalanced in the experiment.
Design
Data Analysis
To assess statistical significance, we utilized Bayes factors (BFs) to determine the extent to which the observed data adjust the a priori belief in the alternative and null hypotheses. Values of 3 << BF << 10 and BF >> 10 indicate moderate and strong evidence against the null hypothesis, respectively. Similarly, values of 1/10 << BF << 1/3 and BF << 1/10 indicate moderate and strong evidence in favor of the null hypothesis, respectively [15, 27, 26].
Estimation of the logistic regression in Equation 2 was performed with Bayesian methods in the JASP (Version 0.19, [14]) environment using the default priors based on the Generalized g-Prior Distribution (CCH; [17]) with $\alpha=0.5$, $\beta=2$, and $s=0$. In addition to Bayes factors for each individual covariate, we also report the 95% credible interval (CI). Although it might be tempting to use the CI to test hypotheses (e.g., rejecting the null hypothesis if the CI does not include the null value), in accordance with recent recommendations [32, 34], we use a more conservative approach, where the CI becomes relevant only after the BF shows evidence for the alternative hypothesis.
Results
@@
We begin our reporting of the results by showing outcomes from objective performance metrics such as the performance of human and collaborative AI agents. Additionally, we highlight behavioral measures related to the degree to which one agent interferes with the plan of the other agent. We then report the results from the subjective metrics based on the questionnaire responses. Finally, we examine human preferences for various types of collaborative agents and apply predictive models to determine which objective and subjective metrics best predict choice.
Objective Metrics
Performance Differences
@@
Figure 5 shows the performance of the human-AI team where the score is combined across the human player and the AI agent. In the high target density condition, human+Omit teams slightly outperformed the human+Ignorant teams, although not statistically significant, $0.4\leq BF_{10}<1$. Thus, the additional design features of the Omit agent presumably caused human performance gains that were at least equally as high as the AI performance decreases. In the low target density condition, the best team performance was achieved with the Ignorant agent closely followed by the Omit agent.
@@
Figure 4: Performance of the human and AI player by AI agent type (columns) and target density (rows) in Experiment 1. Performance is assessed by a relative score: the total points scored relative to the total points that were available during game play. Gray areas visualize the distribution of proportional scores; error bars show the standard error of the mean.
@@
Figure 5: Mean team score by AI agent type and target density in Experiment 1. The team score is based on the sum of score of the AI agent and the human playing with that AI agent relative to the total value of points that was available during game play. Gray shading indicates the distribution of values, while error bars show the standard error from the mean.
Other Behavioral Metrics
@@
We also examined several behavioral metrics that distinguish between the AI agents, focusing on measures related to conflict between the human player and the AI agent. One such metric, the number of AI ’steals’, is defined as the number of times the AI agent intercepts a target initially pursued by the human player. Appendix Figure C2 presents a visualization of these results. As expected, the Ignorant agent shows the highest number of interceptions of targets that the human intended to catch, since this agent disregards the human’s intentions and will pursue targets regardless of the human’s planned actions. Additionally, we analyzed the number of path intersections between the human player and the AI agent as another indicator of potential conflict. Path interceptions were operationalized as the presence of overlap in inter-agent movement trajectories since each agent has an avatar location in the game world and a location they are moving towards. A path intersection thus occurs when the agent’s concurrent trajectories are intersecting.
Subjective Metrics: Questionnaire Responses
@@
Figure 6: Mean questionnaire scores by AI agent type and target density in Experiment 1. Questions were rated on a 7-point Likert-scale. Error bars indicate the standard error from the mean.
@@
Figure 6 shows the questionnaire results. The most general finding that holds for all items except Q3 (“I understood the bot’s intentions”) is that there are significant differences in the ratings for the Ignorant agent, compared to all other agents. These statistical findings reveal a difference between the Ignorant agent and the other agents. However, for some items this pattern is more nuanced as demonstrated by the significant target density interaction effects observed for items Q2, Q5, Q8, $BF_{10}>100$. Exceptions include Q2 (“The bot was competent”) where in the low-density condition the Ignorant, Divide, and Delay agents were rated equally well, while in the high-density condition the Bottom-feeder was rated worse than all other agents, including the Ignorant agent. Furthermore, comparisons in the high-density condition of Q8 (“The bot had a similar playing style to me”) show that the ratings of the Ignorant and Bottom-Feeder agents were roughly equally low, while all other agents received significantly higher ratings. Q3 evinced no differences in ratings across agents.
@@
Table 1 shows example responses when participants were asked to explain their choice. Participants voice many of the human-centered design considerations in their open-ended responses. The theme of teaming was highly represented in our participants’ open-ended responses. Participants frequently pointed out that the Ignorant agent was not being a good teammate. Appendix D provides a content analysis that confirms that the majority of open-ended responses focused on teaming.
Table 1: Examples of explanations provided in the open-ended surveys in Experiment 1. Participants typically referred to the bots using their color labels; for clarity, these references have been replaced with the corresponding bot names.
@@
Teaming “The [Omit] bot felt less like competition and more like a fellow teammate. When I would choose a target, even if it was originally planning on going to that target, it would get out of the way and let me grab the target. This seemed more aligned with two teammates working together than the other robot.” “It felt like we were a team and just trying to collect as many circles as possible where as the [Ignorant] bot felt like it was competing against me and would go change direction based on the highest point circles rather than holding down an area of the platform like me and the [Bottom-Feeder] bot would.”
Likability “The [Divide] bot made playing the game easy and fun. The [Divide] bot spent most of its time in a quadrant away from where I was playing, allowing me not to feel crowded or pressured.”
Intentionality “The [Bottom-Feeder] bot seemed to just let me get any of the targets that I wanted and didn’t try to fight for them.”
Preferences for Collaborative Agents
@@
Figure 7: Choice preferences across pairs of agents for each target density condition in Experiment 1. Each matrix cell indicates the percentage of participants preferring the row-associated agent A over the column-associated agent B. For instance, 82% of participants preferred the Omit agent over the Ignorant agent in the low target density condition. The side column shows the overall preference percentage for each row agent across all pairings. Asterisks denote choice percentages that significantly deviate from 50%, indicated by a Bayes factor greater than 10. Note that results here are averaged across different presentation orders of the agents (e.g., agent A could have been presented first or second in the experiment).
Predictive Models for Human Preferences
@@
To estimate the weights in the logistic regression model of Equation 2, we use Bayesian methods. We separately apply the model to objective and subjective metrics, further breaking down the results by target density. For the objective metrics, we included features such as the human and AI scores, score inequality (defined as the absolute difference between human and AI scores), the number of AI steals, and the number of path intersections between the human and AI agent. For subjective metrics, we included Likert ratings from each of the 8 survey questions. Figure 8 presents the posterior estimates for the $\beta$ weights (see Appendix Table E1 for full results). A positive $\beta$ weight indicates that a larger positive difference in feature values between Agents $X$ and $Y$ increases the likelihood of choosing Agent $X$.
@@
Another key predictor is AI ”steals”, with agents scoring higher on these features being less preferred, although there is only convincing evidence for a non-zero estimate of this effect in the low target density model ($\emph{BF}>100$), perhaps because in low density settings there is more opportunity for competitive interaction.
@@
Figure 8: Posterior ($\beta$) coefficients of Bayesian logistic regression models predicting choice in Experiment 1. The coefficients are shown across objective and subjective metrics (left and right panels) and target densities (indicated by colors). Coefficient estimates can be thought of as weights for the importance of metrics in explaining choices. Error bars represent 95% credible intervals.
Figure 9: Scatter plot of mean relative AI agent score versus the human score in Experiment 1. The results are separated by type of human-AI team and target density. Color is reflective of the probability of the AI type being preferred in pairwise match-ups. The diagonal represents the points of equal scores for the AI agent and the human player. The probability with which an agent is chosen appears inversely related to the distance from the diagonal, suggesting that human players have a preference for score equality. Error bars are indicative of the standard error from the mean.
Trade-off Between Performance and Preference